OpenStudy (anonymous):

Use separation of variables to find the general solution of the differential equation. e^x (y' + 7) = 1 So I'm stumped here, do I move the y'+7 to the other side, or do I distribute the e? I Know I need to separate the x and y, I'm just very confused..

OpenStudy (sdfgsdfgs):

in order to separate the y' from x; try taking ln on both left n right side of the eqn...

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle e^x\left(\frac{dy}{dx}+7\right)=1 }\) \(\large\color{black}{ \displaystyle \left(\frac{dy}{dx}+7\right)=\frac{1}{e^x} }\) \(\large\color{black}{ \displaystyle \frac{dy}{dx}=\frac{1}{e^x}-7 }\) \(\large\color{black}{ \displaystyle \color{red}{\int}\frac{dy}{dx}\color{red}{dx}=\color{red}{\int}\left(\frac{1}{e^x}-7\right)\color{red}{dx} }\)

OpenStudy (solomonzelman):

no need to ln both sides, the left side is simply y+C and the right side you can do.

OpenStudy (sdfgsdfgs):

wait - sorry I misread the eqn. i think @SolomonZelman has the right solution here.

OpenStudy (anonymous):

@SolomonZelman you're a math wiz haha thank you so much, and just to make sure y' is the same as dy/dx right? And thank you @sdfgsdfgs!

OpenStudy (solomonzelman):

well, dy/dx is a more convenient for me notation... also you can visualize how to cancel the differential.

OpenStudy (solomonzelman):

the derivative is the slope after-all, that is: \(\large\color{black}{ \displaystyle \Delta y / \Delta x }\) - and that is \(\large\color{black}{ \displaystyle {\rm d} y / {\rm d} x }\)

OpenStudy (solomonzelman):

well, the notation doesn't matter that much, but I like it that way.

OpenStudy (solomonzelman):

I guess you are done pretty much, bye and yw

OpenStudy (anonymous):

Got it, just making sure, thank you again! :-)

OpenStudy (solomonzelman):

anytime